OpenStudy (anonymous):

Polynomials find all polynomials P(x) with real coefficients such that: \[P(2 P(x))=2 P(P(x))+2 (P(x))^2\]

OpenStudy (asnaseer):

I get:\[P(x)=x^2\]

OpenStudy (anonymous):

yes thats one of the answers

OpenStudy (asnaseer):

Oh - I worked it out as the ONLY answer?

OpenStudy (asnaseer):

P(x) must be a poly of degree 2 at most - correct?

OpenStudy (anonymous):

thats right

OpenStudy (asnaseer):

ok - maybe I made a mistake somewhere in my algebra - I'll try again...

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

BTW: I'm assuming this is a challenge question and not something you need help in? :)

OpenStudy (asnaseer):

ok - gr8 - I do like the questions you put up - they make you think! :)

OpenStudy (asnaseer):

ok, I did make a mistake, so I now get 3 solutions:\[P(x)=x^2+bx\qquad\text{ for some constant }b\]\[P(x)=-\frac{1}{2}\]\[P(x)=0\]

OpenStudy (anonymous):

@asnaseer completely right

OpenStudy (asnaseer):

thanks for letting me know! :)

OpenStudy (anonymous):

yw ;)